How fast would a roc falling from an 85 m high cliff on Mars be traveling when it hit the surface if gMars is 3.75 m/s, using vf-/2gh? m/s
Answer #1

As given in the problem -

g(Mars) = 3.75 m/s^2

Height of the cliff, h = 85.0 m

Also given that -

v(f) = sqrt(2*g*h)--------------------------------------------(i)

You can simply put the value of variables in equation (i) and you will get the value of v(f).

But before this, I want to show you from where equation (i) is derived.

Use the expression -

v2^2 = v1^2 + 2*a*s

Initially, initial velocity, v1 = 0

Acceleration, a = g

s = h

and of course, v2 = vf

So, put the values we get -

vf^2 = 0 + 2*g*h

=> vf = sqrt(2*g*h)

So again put the values as given in the problem -

vf = sqrt(2*3.75*85.0) = 25.25 m/s

Hence, the speed of the rock when it hits the surface = 25.25 m/s